# Incorporating Flexible Image Conditioning into Text-to-Video Diffusion Models without Training

### v1

@@ Abstract
Text-image-to-video (TI2V) generation is a critical problem for controllable video generation using both semantic and visual conditions. Most existing methods typically add visual conditions to text-to-video (T2V) foundation models by finetuning, which is costly in resources and only limited to a few predefined conditioning settings. To tackle this issue, we introduce a unified formulation for TI2V generation with flexible visual conditioning. Furthermore, we propose an innovative training-free approach, dubbed FlexTI2V, that can condition T2V foundation models on an arbitrary amount of images at arbitrary positions. Specifically, we firstly invert the condition images to noisy representation in a latent space. Then, in the denoising process of T2V models, our method uses a novel random patch swapping strategy to incorporate visual features into video representations through local image patches. To balance creativity and fidelity, we use a dynamic control mechanism to adjust the strength of visual conditioning to each video frame. Extensive experiments validate that our method surpasses previous training-free image conditioning methods by a notable margin. We also show more insights of our method by detailed ablation study and analysis.
Figure 1: In this work, we propose FlexTI2V, a novel training-free approach that can add additional image conditioning to off-the-shelf text-to-video foundation models. Then we are able to frame an arbitrary number of images at arbitrary positions in the synthetic video with vivid motion and smooth transitions. In this figure, images with blue edges are condition images, and images with green edges are generated video frames.
@@ 1 Introduction
In recent years, diffusion models [13] have shown excellent performance in text-to-video (T2V) generation [14, 11, 44, 16, 24, 3, 50, 56, 52], serving as video foundation models for various downstream video generation tasks. Given the vagueness and ambiguity of language [26, 25], recent work has investigated synthesizing videos conditioned on RGB images for precise controllability on the appearance of people, objects and background [2, 27, 5], which is as known as image-to-video (I2V) generation. In these models, RGB images can be used as conditions alone [2, 27, 53, 33], or incorporated with textual prompts for accurate guidance [57, 9, 54, 20]. For clarity, we explicitly refer to the latter as text-image-to-video (TI2V) generation following [20, 37].
@@
Figure 2: Comparison with classic TI2V tasks. Our task requires video generation conditioned on any number of images at any positions, which unifies existing classic TI2V tasks. The images with blue and pink edges are condition images, and images with green edges are generated video frames.
@@
To include additional image conditioning in video generation, the most widely-used approach is finetuning pretrained T2V model with paired texts, images and videos. However, the prohibitive demands of high-quality data and computing resources, and the risks of involving unexpected bias from the data during finetuning impede the development of TI2V models. Moreover, as shown in Fig. 2, off-the-shelf TI2V models can only apply the image conditioning setting pre-defined during training (e.g., image animation [9, 58, 4] and video frame interpolation [43, 39]), lacking sufficient flexibility in the number and positions of condition images. Any changes in the conditioning setting (e.g., changes of image numbers or positions) require re-training of the entire model. To avoid the computing overhead and undesired bias of finetuning, training-free image conditioning approaches are promising solutions to incorporate flexible visual guidance into off-the-shelf T2V models.
@@
Prior to our work, recent studies have been conducted on training-free controllable video generation with various conditions [22, 23, 59, 17, 28, 29]. For training-free image conditioning, the previous methods either synthesize only one video frame in one denoising pass, making it inefficient in inference [37], or fill in the missing video frames between two bounded images solely based on visual cues, losing semantic control of generated frames in between [7]. Similar to finetuned TI2V models, they are also only capable of following pre-defined conditioning settings, and still short of flexibility for image conditioning. These inevitable deficiencies make the existing training-free methods suboptimal for TI2V generation problem, thus demanding more efforts and novel solutions.
@@
In this work, we formally define a generic TI2V generation problem with high flexibility in the amounts and positions of condition images (see Figs. 1, 2), unifying the existing TI2V tasks such as image animation, rewinding and video frame interpolation. To address aforementioned issues, we introduce FlexTI2V, an innovative flexible training-free approach that can condition an off-the-shelf text-to-video diffusion model on an arbitrary number of images at arbitrary positions in the synthetic video. Specifically, we first reverse the images to noisy representations at each denoising step. To harmonize the difference in denoising paths of T2V models and image conditioning, we propose a novel and easy-to-implement random patch swapping method to incorporate visual features of input images into each video frame, by swapping a random portion of patches between video frames and condition images. We also use a dynamic adjustment strategy to control the percentage of swapped patches and denoising steps for each frame to prevent frozen videos. The extensive experiments suggest that our approach outperforms the previous training-free image conditioning methods for video diffusion models. Overall, our main contributions are summarized as follows:
  • (1)
@@
  • (2)
    We introduce FlexTI2V, a novel approach to incorporate flexible image conditioning into pretrained text-to-video diffusion models without any requirements of finetuning.
  • (3)
    We propose an innovative random patch swapping method to incorporate visual guidance into the denoising process of video generation in a plug-and-play manner, with a dynamic control strategy to balance creativity and fidelity of video generation.
  • (4)
    The experiments indicate that our approach surpasses the existing training-free image conditioning methods by a notable margin. We further show more insights of our approach via detailed ablation studies and analysis.
@@ 2 Related Work
Text-to-video (T2V) generation has gained impressive advancement since the emergence of diffusion models [14, 11, 15, 3, 24, 56, 50, 52]. Given the ambiguity of language in describing visual details, RGB images are exploited as the condition for video generation, which requires models synthesize smooth video segments with the images exactly as video frames. Image conditioning can be used alone for image-to-video (I2V) generation [1, 6, 53, 33, 2, 27], or integrated with textual prompts for text-image-to-video (TI2V) generation [36, 57, 21, 4, 51, 20]. Here, we mainly discuss previous TI2V methods. Hu et al. [19] propose TI2V task for the first time, and develop a transformer-based architecture to achieve decent controllability and diversity in generated videos. Guo et al. [9] develop a practical framework, AnimateDiff, to train a model-agnostic plug-and-play motion module that can turn any personalized text-to-image diffusion model into an image animator. Xing et al. [54] introduce a novel model – DynamiCrafter, which can animate both human bodies and objects. Fu et al. [8] propose a unified video completion method that can tackle image animation, rewinding and infilling using a single model. Though previous studies have shown promising results for TI2V problem, most methods require training some layers or the entire model to incorporate image conditioning or motion over time. As a result, existing models can only handle TI2V tasks pre-defined in training, such as animation (one image as the first video frame) and infilling (two images as the first and last video frames). In this paper, we introduce a generic TI2V problem that unifies previous TI2V tasks. Our proposed method is able to condition T2V diffusion models on an arbitrary number of images at arbitrary positions of the output videos without finetuning.
Video Frame Interpolation
@@
Video frame interpolation has been extensively studied in recent years, with the model architectures spanning across convolutional networks [38, 31, 39], transformers [8] and diffusion models [7, 40, 41]. The goal of video frame interpolation task is filling the missing video frames between two bounded images with smooth motion and transition. Liu et al. [32] propose a convolutional network to synthesize video frames by flowing pixel values from bounded images. Xu et al. [55] point out that most interpolation methods assume a uniform motion between two consecutive images, thus resulting in oversimplified linear models for interpolation. To generate complex motion flows, They propose a quadratic interpolation method to model acceleration information in videos. Sim et al. [43] collect a large-scale dataset of 4K videos of 1000 fps with extreme motions. They also train XVFI-Net which can model large motions and complex transition of textures between two video frames. Previous work mainly aims at finding a plausible path of transition from the first video frame to the last one. How to go beyond the two bounded images remains understudied. In this work, we investigate the generic TI2V problem including both classic video frame interpolation and frame synthesis beyond the bound of start and end images.
Training-free Approaches for Diffusion Foundation Models
@@
Despite the success of diffusion foundation models in image and video generation, the high computing overhead and large amounts of training data are inevitable bottlenecks when we finetune the foundation models for downstream tasks. Recently, the development of training-free approaches provide an excellent trade-off between performance and cost [12, 35, 48, 30, 23, 22, 37, 47, 7, 59, 17, 28, 29, 18]. Hertz et al. [12] introduce prompt-to-prompt method to edit synthetic images by manipulating self-attention maps, which is further extended to real photo editing using null-text inversion [35]. Kara et al. [22] adapt a pretrained text-guided image editing model for video editing tasks by random noise shuffling. In terms of image-conditioned video generation, the work most related to our method is TI2V-Zero [37]. They duplicate the condition image and synthesize only one video frame in a single denoising process, which enforces the pretrained temporal layer to leverage visual features from the image. However, this method suffers from slow inference speed and compounding error. Feng et al. [7] convert an image-to-video model to a bounded-image interpolation model by merging the two separate denoising paths conditioned on the bounded images. Hou et al. [17] propose to map a single image to 3D point cloud and then synthesize new videos from other viewpoints. Existing training-free TI2V methods usually focus on a specific problem setting, such as image animation and interpolation. In this work, we make the first attempt to develop a training-free method for various image conditioning settings, with high flexibility in image amounts and positions.
@@ 3 Preliminaries
Latent Diffusion Models
Denoising Diffusion Probabilistic Models (DDPM) [13] are initially proposed in RGB space, which is computationally expensive because of the high resolution. As a solution, latent diffusion models (LDM) [42] project RGB images or videos into a low-dimension latent space with a visual encoder $\mathcal{E}$. The denoising process is conducted in the latent space to generate the latent representation of desired images or videos, which are decoded to RGB space by a decoder $\mathcal{D}$. To achieve controllable generation, the conditions are also embedded into the same space, and typically merged into the model through cross-attention layers.
Denoising Diffusion Implicit Models (DDIM)
DDIM [45] basically include two stages – diffusion stage and denoising stage. Given a clean representation $\bm{z}_{0}$ in the latent space, Gaussian noise is added to the $\bm{z}_{0}$ step by step in the diffusion stage. The single-step sample $\bm{z}_{t}$ at step $t$ also follows a Gaussian distribution, $\bm{z}_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\;\bm{z}_{0},\sqrt{1-\bar{\alpha}_{t}}\;\bm{I})$. The closed-form equation is
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. $\beta_{t}$ is the variance schedule. In denoising stage, DDPM build the denoising process as a Markov chain, and estimates the distribution of $\bm{z}_{t-1}$ from $\bm{z}_{t}$. The denoising process is slow due to a large amount of steps (typically 1000 steps). To accelerate sampling, DDIM define a more general non-Markovian process to estimate the Gaussian distribution of $\bm{z}_{t-1}$ conditioned on $\bm{z}_{t}$ and $\bm{z}_{0}$, thus reducing sampling steps significantly. After substitution of $\bm{z}_{0}$ with an expression of $\bm{z}_{t}$, the DDIM sampling is formulated as
$\bm{z}_{t-1}=\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)+\sigma_{t}\cdot\bm{\epsilon}^{\prime},$ (2)
where $\bm{\epsilon}^{\prime}\sim\mathcal{N}(\bm{0},\bm{I})$. $\bm{\varepsilon}_{\theta}$ is a neural network trained to predict the noise added to $\bm{z}_{t}$. $\sigma_{t}$ is a parameter that controls the stochasticity of the sampling process. Setting $\sigma_{t}=0$ yields a deterministic path that can reduce sampling steps with minor loss of generation quality [45]. Then Eq. 2 can be simplified to a function of $\bm{z}_{t}$ and $t$, i.e., $\bm{z}_{t-1}=\bm{\mu}_{\theta}(\bm{z}_{t},t)$.
Figure 3: Overview of the proposed FlexTI2V approach. We invert the condition image embedding to noisy representation $\tilde{\bm{x}}_{t}$ at each step. The final noise $\tilde{\bm{x}}_{T}$ is reused as initialization for video synthesis. At step $t$, we directly replace the video frames with images at the desired positions. Then, for each video frame, we randomly swap a portion of patches with bounded condition images based on the relative distance between the frame and images. Though we show a special case of using two condition images in this figure, our method can naturally extend to any number of images at any positions. Note that all operations of our method occur in the latent space. We visualize RGB images and frames on the latent representations only for intuitive understanding.

### v2

# Unified Text-Image-to-Video Generation: A Training-Free Approach to Flexible Visual Conditioning

@@ Abstract
Text-image-to-video (TI2V) generation is a critical problem for controllable video generation using both semantic and visual conditions. Most existing methods typically add visual conditions to text-to-video (T2V) foundation models by finetuning, which is costly in resources and only limited to a few pre-defined conditioning settings. To tackle these constraints, we introduce a unified formulation for TI2V generation with flexible visual conditioning. Furthermore, we propose an innovative training-free approach, dubbed FlexTI2V, that can condition T2V foundation models on an arbitrary amount of images at arbitrary positions. Specifically, we firstly invert the condition images to noisy representation in a latent space. Then, in the denoising process of T2V models, our method uses a novel random patch swapping strategy to incorporate visual features into video representations through local image patches. To balance creativity and fidelity, we use a dynamic control mechanism to adjust the strength of visual conditioning to each video frame. Extensive experiments validate that our method surpasses previous training-free image conditioning methods by a notable margin. Our method can also generalize to both UNet-based and transformer-based architectures.
Figure 1: We propose FlexTI2V, a novel training-free approach that can add flexible image conditioning to off-the-shelf text-to-video foundation models. We are able to frame an arbitrary number of images at arbitrary positions in the synthetic video with vivid motion and smooth transitions. In this figure, images with blue edges are condition images, and images with green edges are generated video frames.
@@ 1 Introduction
In recent years, diffusion models [ho2020denoising] have shown excellent performance in text-to-video (T2V) generation [ho2022video, he2022latent, singer2023makeavideo, hongcog2023video, kong2024hunyuanvideo, blattmann2023align, wang2025wan, yang2025cogvideox, wang2023modelscope], serving as video foundation models for various downstream video generation tasks. Given the vagueness and ambiguity of language [lai2025unleashing, lai2024lego], recent work has investigated synthesizing videos conditioned on RGB images for precise controllability on the appearance of people, objects and background [blattmann2023stable, li2024generative, chen2024seine], which is as known as image-to-video (I2V) generation. In these models, RGB images can be used as conditions alone [blattmann2023stable, li2024generative, wang2022latent, mahapatra2022controllable], or incorporated with textual prompts for accurate guidance [zhang2023i2vgen, guo2023animatediff, xing2024dynamicrafter, huang2025step]. For clarity, we explicitly refer to the latter as text-image-to-video (TI2V) generation following [huang2025step, ni2024ti2v].
@@
Figure 2: Comparison with classic TI2V tasks. Our task requires video generation conditioned on any number of images at any positions in the output video, which unifies existing classic TI2V tasks. The images with blue and pink edges are condition images, and images with green edges are generated video frames.
@@
To include extra image conditioning in video generation, the most widely-used approach is finetuning pretrained T2V model with paired texts, images and videos. However, the prohibitive demands of high-quality data and computing resources, and the risks of involving unexpected bias from the data during finetuning impede the development of TI2V models. Moreover, as shown in Fig. 2, off-the-shelf TI2V models can only apply the image conditioning setting pre-defined during training (e.g., image animation [guo2023animatediff, zhang2024pia, chen2024livephoto] and video frame interpolation [sim2021xvfi, reda2022film]), lacking sufficient flexibility in the number and positions of condition images. Any changes in the conditioning setting require re-training of the entire model. To avoid the computing overhead and undesired bias of finetuning, training-free image conditioning approaches are promising solutions to incorporate flexible visual control into off-the-shelf T2V models.
@@
Prior to our work, recent studies have been conducted on training-free controllable video generation with various conditions [kara2024rave, khachatryan2023text2video, zhou2024storydiffusion, hou2025training, liang2025looking, liang2025movie]. For training-free image conditioning, the previous methods either synthesize only one video frame in one denoising pass, making it inefficient in inference [ni2024ti2v], or fill in the missing video frames between two bounded images solely based on visual cues, losing semantic control of generated frames in between [feng2024explorative]. Similar to finetuned TI2V models, they are also only capable of following pre-defined conditioning settings, and still short of flexibility for image conditioning. These inevitable deficiencies make the existing training-free methods suboptimal for TI2V generation problem, thus demanding more efforts and novel solutions.
@@
In this work, we formally define a generic TI2V generation problem with high flexibility in the amounts and positions of condition images (see Figs. 1, 2), unifying the existing TI2V tasks such as image animation, rewinding and video frame interpolation. To address aforementioned issues, we introduce FlexTI2V, an innovative flexible training-free approach that can condition an off-the-shelf text-to-video diffusion model on an arbitrary number of images at arbitrary positions in the synthetic video. Specifically, we first reverse the images to noisy representations at each denoising step. To harmonize the difference in denoising paths of T2V models and image conditioning, we propose a novel random patch swapping method to incorporate visual features of input images into each video frame, by swapping a random portion of patches between video frames and condition images. We also use a dynamic adjustment strategy to control the percentage of swapped patches and denoising steps for each frame to prevent frozen videos. The extensive experiments suggest that our approach outperforms the previous training-free image conditioning methods for video diffusion models. Overall, our main contributions are summarized as follows:
@@
  • We introduce FlexTI2V, a novel training-free approach to incorporate flexible image conditioning into pretrained text-to-video diffusion models. The key innovation is a random patch swapping method to incorporate visual guidance into the denoising process of video generation in a plug-and-play manner.
  • The experiments indicate that our approach notably surpasses the existing training-free image conditioning methods. We further validate the generalization of our method to both UNet and transformer architectures.
@@ 2 Related Work
Video Generation Conditioned on Texts and Images Text-to-video (T2V) generation has gained impressive advancement since the emergence of diffusion models [ho2022video, he2022latent, hong2022cogvideo, blattmann2023align, kong2024hunyuanvideo, yang2025cogvideox, wang2025wan, wang2023modelscope]. Given the ambiguity of language in describing visual details, RGB images are exploited as the condition for video generation. Image conditioning can be used alone for image-to-video (I2V) generation [blattmann2021understanding, dorkenwald2021stochastic, wang2022latent, mahapatra2022controllable, blattmann2023stable, li2024generative], or integrated with textual prompts for text-image-to-video (TI2V) generation [ni2023conditional, zhang2023i2vgen, jiang2024videobooth, chen2024livephoto, wang2025dreamvideo, huang2025step]. Here, we mainly discuss previous TI2V methods. hu2022make propose TI2V task for the first time, and develop a transformer-based architecture to achieve decent controllability and diversity in generated videos. guo2023animatediff develop a practical framework, AnimateDiff, to train a model-agnostic plug-and-play motion module that can turn any personalized text-to-image diffusion model into an image animator. xing2024dynamicrafter introduce a novel model – DynamiCrafter, which can animate both human bodies and objects. fu2023tell propose a unified video completion method that can tackle image animation, rewinding and infilling using a single model. Most prior methods require training some layers or the entire model to incorporate image conditioning or motion over time. As a result, existing models can only handle TI2V tasks pre-defined in training, such as animation and infilling. In this paper, we introduce a generic TI2V problem that unifies previous TI2V tasks. Our proposed method is able to condition T2V diffusion models on an arbitrary number of images at arbitrary positions of the output videos without finetuning.
@@
Video Frame Interpolation Video frame interpolation has been extensively studied in recent years, with the model architectures spanning across convolutional networks [niklaus2017video, liu2020enhanced, reda2022film], transformers [fu2023tell] and diffusion models [feng2024explorative, github2024cogvideox, github2025wan]. liu2017video propose a convolutional network to synthesize video frames by flowing pixel values from bounded images. xu2019quadratic point out that most interpolation methods assume a uniform motion between two consecutive images, thus resulting in oversimplified linear models for interpolation. To generate complex motion flows, They propose a quadratic interpolation method to model acceleration information in videos. Previous work mainly aims at finding a plausible path of transition from the first video frame to the last one. How to go beyond the two bounded images remains understudied. In this work, we work on the generic TI2V problem including both classic video frame interpolation and frame synthesis beyond the bound of start and end images.
@@
Training-free Methods for Diffusion Foundation Models The high computing overhead and large amounts of training data of finetune the foundation models are inevitable bottlenecks for downstream tasks. Recently, the development of training-free approaches provide an excellent trade-off between performance and cost [hertz2022prompt, mokady2023null, tumanyan2023plug, liu2024towards, khachatryan2023text2video, kara2024rave, ni2024ti2v, titov2024guide, feng2024explorative, zhou2024storydiffusion, hou2025training, liang2025looking, liang2025movie, hsiao2025tf]. hertz2022prompt introduce prompt-to-prompt method to edit synthetic images by manipulating self-attention maps, which is further extended to real photo editing using null-text inversion [mokady2023null]. kara2024rave adapt a pretrained text-guided image editing model for video editing tasks by random noise shuffling. In terms of image-conditioned video generation, the work most related to our method is TI2V-Zero [ni2024ti2v]. They duplicate the condition image and synthesize only one video frame in a single denoising process, which enforces the pretrained temporal layer to leverage visual features from the image. However, this method suffers from slow inference speed and compounding error. feng2024explorative convert an image-to-video model to a bounded-image interpolation model by merging the two separate denoising paths conditioned on the bounded images. hou2025training propose to map a single image to 3D point cloud and then synthesize new videos from other viewpoints. Existing training-free TI2V methods usually focus on a specific problem setting. In this work, we make the first attempt to develop a training-free method for various image conditioning settings, with high flexibility in image amounts and positions.
@@ 3 Preliminaries
Latent Diffusion Models Denoising Diffusion Probabilistic Models (DDPM) [ho2020denoising] are initially proposed in RGB space, which is computationally expensive because of the high resolution. As a solution, latent diffusion models (LDM) [rombach2022high] project RGB images or videos into a low-dimensional latent space with a visual encoder $\mathcal{E}$. The denoising process is conducted in the latent space to generate the latent representation of desired images or videos, which are decoded to RGB space by a decoder $\mathcal{D}$. To achieve controllable generation, the conditions are also embedded into the same space, and typically merged into the model through cross-attention layers.
Figure 3: Overview of the proposed FlexTI2V approach. We invert the condition image embedding to noisy representation $\tilde{\bm{x}}_{t}$ at each step. The final noise $\tilde{\bm{x}}_{T}$ is reused as initialization for video synthesis. At step $t$, we directly replace the video frames with images at the desired positions. Then, for each video frame, we randomly swap a portion of patches with bounded condition images based on the relative distance between the frame and images. Though we show a special case of using two condition images in this figure, our method can naturally extend to any number of images at any positions. Note that all operations of our method occur in the latent space. We visualize RGB images and frames on the latent representations simply for intuitive understanding.
Denoising Diffusion Implicit Models (DDIM) DDIM [song2020denoising] basically include two stages – diffusion stage and denoising stage. Given a clean representation $\bm{z}_{0}$ in the latent space, Gaussian noise is added to the $\bm{z}_{0}$ step by step in the diffusion stage. The single‑step sample $\bm{z}_{t}$ at step $t$ also follows a Gaussian distribution, $\bm{z}_{t}\sim\mathcal{N}(\sqrt{\bar{\alpha}_{t}}\;\bm{z}_{0},\ (1-\bar{\alpha}_{t})\bm{I})$. The closed-form equation is
@@
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ and $\alpha_{t}=1-\beta_{t}$. $\beta_{t}$ is the noise schedule. In denoising stage, DDPM build the denoising process as a Markov chain, and estimates the distribution of $\bm{z}_{t-1}$ from $\bm{z}_{t}$. To accelerate sampling, DDIM define a more general non-Markovian process to estimate the Gaussian distribution of $\bm{z}_{t-1}$ conditioned on $\bm{z}_{t}$ and $\bm{z}_{0}$, thus reducing sampling steps significantly. After substitution of $\bm{z}_{0}$ with an expression of $\bm{z}_{t}$, DDIM sampling is formulated as
$\bm{z}_{t-1}=$ $\frac{\sqrt{\bar{\alpha}_{t-1}}}{\sqrt{\bar{\alpha}_{t}}}\left(\bm{z}_{t}-\sqrt{1-\bar{\alpha}_{t}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)\right)$ (2)
$+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\bm{\varepsilon}_{\theta}(\bm{z}_{t},t)+\sigma_{t}\cdot\bm{\epsilon}^{\prime}$
where $\bm{\epsilon}^{\prime}\sim\mathcal{N}(\bm{0},\bm{I})$. $\bm{\varepsilon}_{\theta}$ is a neural network trained to predict the noise added to $\bm{z}_{t}$. $\sigma_{t}$ is a parameter that controls the stochasticity of the sampling process. Setting $\sigma_{t}=0$ yields a deterministic path that can reduce sampling steps with minor loss of generation quality [song2020denoising]. Then Eq. 2 can be simplified to a function of $\bm{z}_{t}$ and $t$, i.e., $\bm{z}_{t-1}=\bm{\mu}_{\theta}(\bm{z}_{t},t)$.